OpenStudy (anonymous):

what is the solution set for 5x-9=11x+15 ?

OpenStudy (anonymous):

get the x's on one side and get the 9 and 15 one one side.

OpenStudy (anonymous):

x=-4

OpenStudy (anonymous):

oh okay so how would you solve it?

OpenStudy (anonymous):

like the steps please

OpenStudy (anonymous):

get all the ones with x on one side by adding or subtracting. so you have 5x-9=11x+15. Then you would do 5x-9(+9)=11x+15 so then you have 5x=11x+15+9. then you do the same with the 11x. subtract 11x from both sides so that you have 5x-11x=15+9. You get -6x=24. divide the -6 by both sides. so then you have -6x/-6=24/-6. the two 6's cancel out so then you have x=-4.

OpenStudy (anonymous):

thank you@chelseachels (:

OpenStudy (anonymous):

your welcome